OpenStudy (anonymous):

Solve. 2(p + 1) = 16 (Points : 4) 7 7.5 8.5 9

OpenStudy (anonymous):

7

OpenStudy (kamibug):

1. Distribute 2p + 2 = 16 2. Subtract both sides by 2 2p = 14 3. Divide both sides by 2. p = 14 / 2 = ?

OpenStudy (anonymous):

2(p + 1) = 16 You first have to use the distributive property. So, multiply 2 by everything in the parentheses. 2p + 2 = 16 Then, subtract 2 from both sides to isolate the variable to one side. 2p = 14 Then, just divide two from both sides. p = 7. So, the answer is the first one. (: